OpenStudy (anonymous):

How many solutions are there to the equation 14x+2=12x?

OpenStudy (anonymous):

one

OpenStudy (austinl):

Solve for x.

OpenStudy (anonymous):

14X+2=12X =>14X-12X=-2 =>2X=-2 =>X=-1

OpenStudy (austinl):

14x+2=12x 2x+2=0 2x=-2 x=-1

OpenStudy (chmvijay):

only one :)

OpenStudy (anonymous):

Thanxx